A gas is heated from 263.0 K to 298.0 K and the volume is increased from 24.0 liters to 35.0 liters by moving a large piston within a cylinder. If the original pressure were 1.00 atm, what would the final pressure be? 


1
Expert's answer
2021-05-17T04:08:27-0400

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


"\frac{1.0×24.0}{263.0}=\frac{P_2×35.0}{298.0}"


"P_2=0.777atm"